9. Let X have an exponential distribution with A 1 (see Question 5), and let Y log(X). Find the probability density function of Y. Where is the density non-zero? Note that in this course, log refers to the log base e, or natural log, often symbolized In. The distribution of Y is called the (standard) Gumbel, or extreme value distribution.
Answer #1

9)

f(x) = e^(-x)

F(x) = 1 - e^(-x)

P(Y < y)

= P(log x < y)

= P(x < e^y)

= F(e^y)

= 1 - e^(-e^y)

pdf = d/dy F(y)

day (